the perimeter of a parallelogram is 54cm. the ratio of the lengths of the sides is 2:7. what are the lengths of the sides?
Answer by user_dude2008(1862)

2/7=a/b
2b=7a

2a+2b=54
2a+7a=54
9a=54
a=6

2b=7*6
2b=42
b=21

sides are 6, 21